A balloon is floating around outside your window. The temperature outside is -7 ∘C , and the air pressure is 0.800 atm . Your neighbor, who released the balloon, tells you that he filled it with 3.20 moles of gas. What is the volume of gas inside this balloon? Express your answer to three significant figures and include the appropriate units.
Answer #1

Given that

Temperature (T) = -7oC = -7 + 273 = 266 K

Pressure (P) = 0.800 atm

Moles of gas (n) = 3.20 mol

Gas constant (R) = 0.0821 L. atm / mol.K

Let V be the volume of gas

From Ideal gas equation

PV = nRT

V = nRT/P = (3.20 mol * 0.0821 L. atm / mol.K * 266K) / 0.800 atm = 87.4 L